# Enabling Physical Localization of Uncooperative Cellular Devices

### v2

@@ 7 Discussion
While we utilized multiangulation-based localization, the proposed approaches can also benefit multilateration-based localization. For instance, we can improve LTrack [31] utilizing ToA-based multilateration, by addressing challenges such as tracking RNTI associated with online identity instead of IMSI extraction, ensuring continuous uplink transmission during multiple ToA measurements, and achieving ToA in the case with low signal power. Furthermore, attackers employing LTrack can overcome the impact of repeaters by differentiating the signal through power boosting.
Thanks to UMA, they can achieve reliable fine-grained localization in realistic scenarios such as vishing investigation.
Multipath effects. The proposed attack considers realistic scenarios, yet additional challenges might arise in multipath environments due to frequent reflections caused by obstructions. Fortunately, existing research has explored strategies for radio direction detection [49, 52] and indoor localization [14, 39, 55, 32]. Thus, we believe integrating those algorithms can make localization accurate in such environments.
Countermeasure. First, one plausible method to mitigate a power-boosting attack is to encrypt the radio traffic carrying DCI 0. Second, as a potential solution to prevent scheduling manipulation attack, allocating the radio configuration for scheduling request over an encrypted channel can be considered. One possible approach is to reallocate user-dedicated radio configuration, using RRC Connection Reconfiguration. By doing so, it becomes challenging for the attacker to identify the victim’s schedulingRequestConfig, which is crucial for manipulating uplink scheduling. While UMA is a permanent and unpatchable attack due to the finalization of the LTE specification, ongoing 5G/6G projects can potentially mitigate the threat posed by UMA with these countermeasures.
8 Conclusion
@@
In this paper, we present UMA, enabling the physical localization of uncooperative cellular devices. Our work shows that a localizer with only the target’s online identity can determine the physical location of the target by actively exploiting vulnerabilities in LTE network protocols. Through ensuring strict compliance with legal boundaries, we have demonstrated the attack feasibility and effectiveness of UMA. We emphasize that UMA is a permanent and unpatchable attack since it utilizes vulnerabilities of LTE networks of which specification development is complete.
Appendix A Acronyms
3GPP Third Generation Partnership Project
AoA Angle of Arrival
BSR Buffer Status Report
CDF Cumulative Density Function
CID Cell ID
COTS Commercial off-the-shelf
CVD Coordinated Vulnerability Disclosure
C-RNTI Cell RNTI
DCI Downlink Control Information
DFing Direction Finding
DL DownLink
DRB Data Radio Bearer
eNB Evolved Node B
E-CID Enhanced Cell ID
GSMA Global System for Mobile Communications Association
GUTI Globally Unique Temporary Identity
LCID Logical Channel ID
LTE Long Term Evolution
MAC Medium Access Control
MIB Master Information Block
MUSIC Multiple Signal Classification
OTDoA Observed Time Difference of Arrival
PDCCH Physical Downlink Control CHannel
PDSCH Physical Downlink Shared Channel
PHICH Physical HybridARQ Indicator Channel
PHY Physical
PSS Primary Synchronization Signal
PUCCH Physical Uplink Control Channel
PUSCH Physical Uplink Shared Channel
QoS Quality of Service
RF Radio Frequency
RNTI Radio Network Temporary Identifier
RRC Radio Resource Control
SDR Software Defined Radio
SMS Short Message Service
SRB Signaling Radio Bearer
SSS Secondary Synchronization Signal
TA Timing Advance
TDoA Time Difference of Arrival
TMSI Temporary Mobile Subscriber Identity
ToA Time of Arrival
TPC Transmit Power Control
Tx PWR Transmission Power
UE User Equipment
UL UpLink
UMA Uncooperative Multiangulation Attack
UTDoA Uplink Time Difference of Arrival
Appendix B Localization Attack in the Wild
B.1 LTE Channel Sniffer
Figure 9: LTE channel sniffer.
To get hands-on experience with the localization attack with multiangulation, we first build the sniffer. Then, to examine the performance of the attack in the wild, we conduct localization in the commercial cellular network.
Note that we do not use specialized hardware such as LTEye [32]. Instead, we use a parabolic antenna [11] to eavesdrop on RF signals of downlink and uplink channels. The resulting LTE channel sniffer runs on an Intel Core i5 desktop with SDR. As shown in  Figure 9, the left channel of the USRP is used for the downlink channel eavesdropping, and the right channel is for the uplink channel eavesdropping. A directional antenna is mounted on the uplink channel of the USRP, and both channels share the same clock.
B.2 Experiment of Localization Attack
Figure 10: Estimated location of the UE using three sniffers.
(a) $\theta$ Accuracy
(b) Distance Accuracy
Figure 11: Angular and distance error of estimated location.
Using the LTE channel sniffer, we conduct localization with three sniffers and use a COTS device (i.e., Galaxy S10) as a target. The sniffers and the UE are deployed in the same cellular coverage. The experiment is performed in a Line-of-Sight (LoS) environment of $100m\times 100m$ size and we deploy three sniffers, as denoted by the purple square in Figure 10. The UE is deployed at 30m intervals in nine places, shown as red circles in Figure 10.
First, we estimate the direction of the strongest uplink signal detected at each sniffer while rotating the directional antenna (i.e., AoA measurement). Then, we perform multiangulation based on AoA measurements of each sniffer to predict the location of the UE. The predictions for the UE’s location at each point are denoted as three yellow crosses in Figure 10. Since we already know the location of the sniffers, we could estimate the location of the target UE in coordinate units.
We evaluate the performance of AoA measurements with a cumulative density function (CDF) of angular error as shown in Figure 11 (a). The maximum angular error of AoA is about ${}^{\circ}$ and the minimum angular error is about ${}^{\circ}$. This might include human error during the process of rotating the antenna and estimating the direction of the strongest uplink signal. Then, we evaluate the performance of location predictions with a CDF of distance error. As shown in Figure 11 (b), the maximum distance error is about 14m and the minimum distance error is about 10m. When two sniffers are used to estimate the location of an UE, it shows better localization performance than the case of combining the measurements of three sniffers.
B.3 Shadow Area: Impact of the UE’s Position
Figure 12: Arrival signal strength according to the distance between the UE and the sniffer.
We measure the arrival signal strength at the sniffer, which runs on a 10Mhz bandwidth LTE channel in an 800Mhz band. We use a COTS device, Samsung Galaxy S10, as the victim’s UE. Both the UE and the sniffer are connected to the commercial network and located within the same commercial cell coverage in an urban environment. In this experiment, we set the parabolic antenna [11] mounted on the sniffer to direct to the UE. To calculate the signal strength, we take an average of the monitored strength at the sniffer over 2,000 subframes during each measurement trial.
First, we consider the case in which the UE is located far from the eNB. We place the UE 90m away from the eNB. We then place the sniffer in the middle of the eNB and the UE. The sniffer measures the arrival signal strength every 10m between the eNB and the UE. Second, we consider the case when the UE is located close to the eNB, in contrast to the first case. For this, we place the UE 10m away from the eNB. The sniffer is situated 20m from the eNB, which is 10m from the UE. Then, we measure the signal strength at the sniffer by placing it every 10m up to 80m from the UE.
In Figure 12, orange-colored and blue-colored lines show the measured signal strength at the sniffer. When the UE is far from the eNB (orange line), the attacker’s sniffer is able to detect the victim’s uplink signals well, up to 80m away from the UE. However, when the UE is close to the eNB (blue line), the sniffer can no longer detect the victim’s uplink signals in the case the sniffer is situated more than 70m away from the UE. This is because the UE transmits its uplink data with lower signal power when the UE is located close to eNB.
B.4 Impact of the UE’s Behavior
Figure 13: Arrival signal strength by the UE’s behavior.
We investigate the impact of a UE’s behavior on the attacker’s sniffer, the UE using the three behaviors employed in the previous experiment (i.e., voice call, web surfing, and bulk uploading). In this experiment, we place the sniffer 45m away from an eNB. We then measure the arrival signal strength by moving the UE from 10m to 150m away from the sniffer at intervals of 10m.
Figure 13 shows the measured signal strength at the attacker’s sniffer over three types of the UE’s behavior. The red box in the figure represents the detection threshold, and it was determined by measuring signal strength when uplink signals do not exist. The attacker’s sniffer easily detects the uplink signals when the UE is doing bulk uploads or making voice calls. On the other hand, the measured signal strength during web surfing is lower than the others, which causes the attacker to fail in detecting the signal when the UE is more than 60m away from the sniffer.
Appendix C Additional Considerations
C.1 Sniffer Deployment
Figure 14: Sniffer deployment in three-sectorized LTE cell.
Here we present a simple strategy in conducting UMA in the real-world environment considering practical issues. One of several sniffers used for multiangulation should be placed nearby an eNB, as a reference point. This is because if the victim is located nearby the eNB, the victim’s UE transmits uplink data with low signal power.
After AoA measurement at the reference point, the attacker can place one or two sniffers at other points based on the estimated direction at the reference point. In general, an LTE cell consists of three sector sites [6], and each sector is in charge of ${}^{\circ}$ range of radio networks (i.e., $3\times 120^{\circ}$). Thus, the attacker can place the sniffers in one sector range as shown in Figure 14. If the attacker tries three-dimensional localization, it is recommended to use three sniffers for multiangulation. Otherwise, using only two sniffers is enough since we have empirically confirmed that it shows better performance for two-dimensional localization.
Finally, when the shadow area occurs during the localization attack, the attacker can enforce the victim’s UE to increase transmission power by exploiting power boosting attack which we discuss in § 4.3.

### v3

@@ 8. Related Work
Through scheduling manipulation, UMA enables LTEye to perform physical localization without its practical requirements; 1) the target UE should be in an active status while the SAR’s mobile antenna rotates 360°and 2) RNTI mapping process with RF fingerprinting is needed when the target’s RNTI is reassigned. While we utilized multiangulation-based localization, UMA can also benefit the ToA-based multilateration technique. For instance, we can improve LTrack (Kotuliak et al., 2022) utilizing multilateration, by addressing challenges such as tracking RNTI associated with online identity instead of IMSI extraction, ensuring continuous uplink transmission during multiple ToA measurements. Lastly, when LTEye and LTrack attackers face challenges stemming from low signal power or the presence of the repeater, they can perform the power boosting attack to improve their localization performance.
Other cellular localization. Lakshmanan et al(Lakshmanan et al., 2021) demonstrated unprivileged location tracking by exploiting Carrier Aggregation (CA) side channels. However, it requires collecting CA measurements across all possible target paths, limiting the capability of localizers when attempting to trace cellular devices. Michalevsky et al(Michalevsky et al., 2015) showed that the location path is leaked by analyzing the power consumption of the target’s device, but it requires installing malware on the device for location tracking. Previous studies (Bitsikas et al., 2023; Baek et al., 2023; Cheng et al., 2023; Hong et al., 2018; Kune et al., 2012; Ludant et al., 2023; Shaik et al., 2016) aimed to determine the presence of the target UE within or beyond a specific cell coverage area by utilizing calls/SMSes. However, these approaches have restricted localization granularity, at best achieving cell-level accuracy.
Message forgery attacks. Tan et al(Tan et al., 2021) proposed data-plane signaling forgeries at unprotected layers in cellular networks. They exploited various vulnerabilities in lower-layer protocols such as BSR and TA to demonstrate threats derived from unprotected control messages. Studies (Erni et al., 2022; Yang et al., 2019; Tan et al., 2021; Kotuliak et al., 2022) have shown that unprotected cellular messages could be overwritten at the PHY layer through SigOver attacks.
9. Conclusion
@@
In this paper, we investigate realistic challenges for the physical localization of uncooperative cellular devices down to the front door. We present UMA as a solution, enabling a localizer with only the target’s online identity (e.g., phone number) to determine the target’s physical location by actively exploiting three vulnerabilities in the LTE specification. While ensuring strict compliance with legal boundaries, we demonstrate the feasibility and effectiveness of UMA. We emphasize that UMA is universally applicable in any LTE network as it utilizes vulnerabilities inherent in cellular specifications. We have responsibly disclosed our findings to GSMA, who confirmed that UMA makes physical localization reliable. Finally, the countermeasures discussed in this paper could be considered for 5G and 6G implementations and specifications.
Acknowledgements.
We sincerely appreciate reviewers and shepherd for their valuable comments. We would also like to thank Mincheol Son, Kwangmin Kim, Jiwoo Seo, Hansung Bae, Taisic Yun, and Beomseok Oh for their assistance in improving our paper. This research was supported and funded by the Korean National Police Agency [Tracking and identifying devices and call traffic in voice phishing ecosystem / PR10-03-020-22].